# Sampling the space of solutions of an artificial neural network

### v1

@@ B.3 Large $K$ limit
$\Psi_{rs}$ $\equiv\langle\langle\varphi_{r}\rangle_{\lambda}\langle\varphi_{s}\rangle_{\lambda}\rangle_{x}-\langle\varphi_{r}\rangle_{\lambda,x}\langle\varphi_{s}\rangle_{\lambda,x}$ (54a)
$\Delta_{r}$ $\equiv\langle\varphi^{2}\rangle_{\lambda,x}-\langle\langle\varphi_{r}\rangle_{\lambda}^{2}\rangle_{x}$ (54b)
$T$ $\equiv\langle\langle\tilde{\varphi}\rangle^{2}_{\lambda}\rangle_{x}-\langle\tilde{\varphi}\rangle_{\lambda,x}^{2}$ (54c)
$\Xi$ $\equiv\langle\tilde{\varphi}^{2}\rangle_{\lambda,x}-\langle\langle\tilde{\varphi}\rangle^{2}_{\lambda}\rangle_{x}$ (54d)
$U_{r}$ $\equiv\langle\langle\varphi_{r}\rangle_{\lambda}\langle\tilde{\varphi}\rangle_{\lambda}\rangle_{x}-\langle\varphi\rangle_{\lambda,x}\langle\tilde{\varphi}\rangle_{\lambda,x}$ (54e)
$\Omega_{r}$ $\equiv\langle\langle\tilde{\varphi}\varphi_{r}\rangle_{\lambda}\rangle_{x}-\langle\langle\tilde{\varphi}\rangle_{\lambda}\langle\varphi_{r}\rangle_{\lambda}\rangle_{x}$ (54f)
@@ Appendix C Effective order parameters
$\varphi_{r}(\lambda_{r},x)$ $\equiv\varphi\left(\sqrt{Q-q_{r}}\lambda_{r}-\sum_{s}\mathcal{T}_{rs}x_{s}\right)$ (70a)
$\tilde{\varphi}(\gamma;\lambda_{1},\lambda_{2},x_{1},x_{2})$ $\equiv\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{r}-\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)$ (70b)
@@
Let’s start by analyzing the terms $\Psi_{rs}\equiv\langle\langle\varphi_{r}\rangle_{\lambda}\langle\varphi_{s}\rangle_{\lambda}\rangle_{x}-\langle\varphi_{r}\rangle_{\lambda,x}\langle\varphi_{s}\rangle_{\lambda,x}$ and $\Delta_{r}\equiv\langle\varphi^{2}\rangle_{\lambda,x}-\langle\langle\varphi_{r}\rangle_{\lambda}^{2}\rangle_{x}$, which only involve $\varphi_{r}$. First notice that
$\langle\varphi_{r}\rangle_{\lambda,x}$ $=\int Dy\,\varphi(\sqrt{Q}y)=\sqrt{\Delta_{Q}(0)}\,,\qquad r=1,2$ (71a)
$\langle\varphi_{r}^{2}\rangle_{\lambda,x}$ $=\int Dy\,\varphi^{2}(\sqrt{Q}y)=\Delta_{Q}(Q)\,,\qquad r=1,2$ (71b)
@@
$\begin{split}\langle\langle\varphi_{1}\rangle_{\lambda}\langle\varphi_{2}\rangle_{\lambda}\rangle_{x}&=\int Dx_{1}Dx_{2}D\lambda_{1}D\lambda_{2}\,\varphi\left(\sqrt{Q-q_{1}}\lambda_{1}+\sum_{s}\mathcal{T}_{1s}x_{s}\right)\,\varphi\left(\sqrt{Q-q_{2}}\lambda_{2}+\sum_{s}\mathcal{T}_{2s}x_{s}\right)\\ &=\int Dx_{1}Dx_{2}D\lambda_{1}D\lambda_{2}\,\varphi\left(\sqrt{Q-q_{1}}\lambda_{1}+\sqrt{q_{1}}x_{1}\right)\,\varphi\left(\sqrt{Q-q_{2}}\lambda_{2}+\frac{p}{\sqrt{q_{1}}}x_{1}+\sqrt{q_{2}-\frac{p^{2}}{q_{1}}}x_{2}\right)\\ &=\int Dx_{1}Dx_{2}\,\varphi\left(\sqrt{Q}x_{1}\right)\,\varphi\left(\frac{p}{\sqrt{Q}}\,x_{1}+\sqrt{Q-\frac{p^{2}}{Q}}x_{2}\right)=\Delta_{Q}(p)\\ \end{split}$ (72)
@@
$\begin{split}\langle\langle\varphi_{r}\rangle_{\lambda}\langle\varphi_{r}\rangle_{\lambda}\rangle_{x}=\langle\langle\varphi_{r}\rangle_{\lambda}^{2}\rangle_{x}=\Delta_{Q}(q_{r})\end{split}$ (73)
@@
Secondly let’s analyze the terms which contain only $\tilde{\varphi}$, i.e. $T\equiv\langle\langle\tilde{\varphi}\rangle^{2}_{\lambda}\rangle_{x}-\langle\tilde{\varphi}\rangle_{\lambda,x}^{2}$ and $\Xi\equiv\langle\tilde{\varphi}^{2}\rangle_{\lambda,x}-\langle\langle\tilde{\varphi}\rangle^{2}_{\lambda}\rangle_{x}$. The terms $\langle\tilde{\varphi}\rangle_{\lambda,x}$ and $\langle\tilde{\varphi}^{2}\rangle_{\lambda,x}$ are easy to analyze since the integrand involves a sum of 4 uncorrelated Gaussian variables, which is Gaussian. We therefore get
$\langle\tilde{\varphi}\rangle_{\lambda,x}$ $=\int Dy\,\varphi(\sqrt{Q}y)=\sqrt{\Delta_{Q}(0)}\,,$ (74a)
$\langle\tilde{\varphi}^{2}\rangle_{\lambda,x}$ $=\int Dy\,\varphi^{2}(\sqrt{Q}y)=\Delta_{Q}(Q)\,.$ (74b)
@@
$\begin{split}\langle\langle\tilde{\varphi}\rangle^{2}_{\lambda}\rangle_{x}&=\int Dx_{1}Dx_{2}\left[\int D\lambda_{1}D\lambda_{2}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{r}+\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)\right]^{2}\\ &=\int Dx\left[\int Dy\,\varphi\left(\frac{\sqrt{Q-2Q\gamma(1-\gamma)-\gamma^{2}q_{1}-(1-\gamma)^{2}q_{2}}}{c_{\gamma}}y+\frac{\sqrt{\gamma^{2}q_{1}+(1-\gamma)^{2}q_{2}+2\gamma(1-\gamma)p}}{c_{\gamma}}\,x\right)\right]^{2}\\ &=\Delta_{Q}\left(\frac{\gamma^{2}q_{1}+(1-\gamma)^{2}q_{2}+2\gamma(1-\gamma)p}{c_{\gamma}^{2}}\right)\end{split}$ (75)
@@
The last computation concerns the correlations between the function $\varphi_{r}$ and $\tilde{\varphi}$, which appear in the variables $U_{r}\equiv\langle\langle\varphi_{r}\rangle_{\lambda}\langle\tilde{\varphi}\rangle_{\lambda}\rangle_{x}-\langle\varphi\rangle_{\lambda,x}\langle\tilde{\varphi}\rangle_{\lambda,x}$, $\Omega_{r}\equiv\langle\langle\tilde{\varphi}\varphi_{r}\rangle_{\lambda}\rangle_{x}-\langle\langle\tilde{\varphi}\rangle_{\lambda}\langle\varphi_{r}\rangle_{\lambda}\rangle_{x}$. We need therefore to evaluate the following two quantities $\langle\langle\tilde{\varphi}\varphi_{r}\rangle_{\lambda}\rangle_{x}$ and $\langle\langle\tilde{\varphi}\rangle_{\lambda}\langle\varphi_{r}\rangle_{\lambda}\rangle_{x}$, for $r=1,2$. We are going to analyze the case $r=1$, the other can be obtained by symmetry. We have
@@
$\begin{split}\langle\langle\tilde{\varphi}\varphi_{1}\rangle_{\lambda}\rangle_{x}&=\int Dx_{1}Dx_{2}D\lambda_{1}D\lambda_{2}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{r}+\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)\varphi\left(\sqrt{Q-q_{1}}\lambda_{1}+\sum_{s}\mathcal{T}_{1s}x_{s}\right)\\ &=\int Dx_{1}Dx_{2}D\lambda_{1}D\lambda_{2}\,\varphi(x_{1})\\ &\times\varphi\left(\frac{\gamma Q+(1-\gamma)p}{\sqrt{Q}c_{\gamma}}x_{1}-\frac{(1-\gamma)p\sqrt{Q-q_{1}}}{\sqrt{q_{1}Q}c_{\gamma}}\lambda_{1}+\frac{(1-\gamma)\sqrt{Q-q_{2}}}{c_{\gamma}}\lambda_{2}+\frac{1-\gamma}{c_{\gamma}}\sqrt{q_{2}-\frac{p^{2}}{q_{1}}}x_{2}\right)\\ &=\int Dx_{1}Dx_{2}\,\varphi(\sqrt{Q}x_{1})\varphi\left(\frac{\gamma Q+(1-\gamma)p}{\sqrt{Q}c_{\gamma}}x_{1}+\frac{1-\gamma}{c_{\gamma}}\sqrt{Q-\frac{p^{2}}{Q}}x_{2}\right)=\Delta_{Q}\left(\frac{\gamma Q+(1-\gamma)p}{c_{\gamma}}\right)\end{split}$ (76)
@@
Notice that this does not depend on either $q_{1}$ nor $q_{2}$. The case $r=2$ can be obtained by sending $\gamma\to 1-\gamma$ and $q_{1},q_{2}\to q_{1},q_{2}$. Let’s analyze the term $\langle\langle\tilde{\varphi}\rangle_{\lambda}\langle\varphi_{r}\rangle_{\lambda}\rangle_{x}$ in the $r=1$ case
@@
$\begin{split}\langle\langle\tilde{\varphi}\rangle_{\lambda}\langle\varphi_{1}\rangle_{\lambda}\rangle_{x}&=\int Dx_{1}Dx_{2}D\lambda_{1}D\lambda_{2}D\lambda_{3}\,\varphi\left(\sqrt{Q-q_{1}}\lambda_{3}+\sum_{s}\mathcal{T}_{1s}x_{s}\right)\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{r}+\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)\\ &=\int Dx_{1}Dx_{2}\,\varphi(\sqrt{Q}x_{1})\varphi\left(\frac{\gamma q_{1}+(1-\gamma)p}{\sqrt{Q}c_{\gamma}}x_{1}+\sqrt{Q-\frac{(\gamma q_{1}+(1-\gamma)p)^{2}}{Qc_{\gamma}^{2}}}x_{2}\right)\\ &=\Delta_{Q}\left(\frac{\gamma q_{1}+(1-\gamma)p}{c_{\gamma}}\right)\end{split}$ (77)
@@
$\begin{split}\langle\langle\tilde{\varphi}\rangle_{\lambda}\langle\varphi_{2}\rangle_{\lambda}\rangle_{x}=\Delta_{Q}\left(\frac{(1-\gamma)q_{2}+\gamma p}{c_{\gamma}}\right)\end{split}$ (78)

### v2

@@ B.3 Large $K$ limit
$\Psi_{rs}$ $\equiv\langle\langle\varphi_{r}\rangle_{v}\langle\varphi_{s}\rangle_{v}\rangle_{x}-\langle\varphi_{r}\rangle_{v,x}\langle\varphi_{s}\rangle_{v,x}$ (56a)
$\Delta_{r}$ $\equiv\langle\varphi^{2}\rangle_{v,x}-\langle\langle\varphi_{r}\rangle_{v}^{2}\rangle_{x}$ (56b)
$T$ $\equiv\langle\langle\tilde{\varphi}\rangle^{2}_{v}\rangle_{x}-\langle\tilde{\varphi}\rangle_{v,x}^{2}$ (56c)
$\Xi$ $\equiv\langle\tilde{\varphi}^{2}\rangle_{v,x}-\langle\langle\tilde{\varphi}\rangle^{2}_{v}\rangle_{x}$ (56d)
$U_{r}$ $\equiv\langle\langle\varphi_{r}\rangle_{v}\langle\tilde{\varphi}\rangle_{v}\rangle_{x}-\langle\varphi\rangle_{v,x}\langle\tilde{\varphi}\rangle_{v,x}$ (56e)
$\Omega_{r}$ $\equiv\langle\langle\tilde{\varphi}\varphi_{r}\rangle_{v}\rangle_{x}-\langle\langle\tilde{\varphi}\rangle_{v}\langle\varphi_{r}\rangle_{v}\rangle_{x}$ (56f)
@@ Appendix C Effective order parameters
$\varphi_{r}(v_{r},x)$ $\equiv\varphi\left(\sqrt{Q-q_{r}}v_{r}-\sum_{s}\mathcal{T}_{rs}x_{s}\right)$ (72a)
$\tilde{\varphi}(\gamma;v_{1},v_{2},x_{1},x_{2})$ $\equiv\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{r}-\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)$ (72b)
@@
Let’s start by analyzing the terms $\Psi_{rs}\equiv\langle\langle\varphi_{r}\rangle_{v}\langle\varphi_{s}\rangle_{v}\rangle_{x}-\langle\varphi_{r}\rangle_{v,x}\langle\varphi_{s}\rangle_{v,x}$ and $\Delta_{r}\equiv\langle\varphi^{2}\rangle_{v,x}-\langle\langle\varphi_{r}\rangle_{v}^{2}\rangle_{x}$, which only involve $\varphi_{r}$. First notice that
$\langle\varphi_{r}\rangle_{v,x}$ $=\int Dy\,\varphi(\sqrt{Q}y)=\sqrt{\Delta_{Q}(0)}\,,\qquad r=1,2$ (73a)
$\langle\varphi_{r}^{2}\rangle_{v,x}$ $=\int Dy\,\varphi^{2}(\sqrt{Q}y)=\Delta_{Q}(Q)\,,\qquad r=1,2$ (73b)
@@
$\begin{split}\langle\langle\varphi_{1}\rangle_{v}\langle\varphi_{2}\rangle_{v}\rangle_{x}&=\int Dx_{1}Dx_{2}Dv_{1}Dv_{2}\,\varphi\left(\sqrt{Q-q_{1}}v_{1}+\sum_{s}\mathcal{T}_{1s}x_{s}\right)\,\varphi\left(\sqrt{Q-q_{2}}v_{2}+\sum_{s}\mathcal{T}_{2s}x_{s}\right)\\ &=\int Dx_{1}Dx_{2}Dv_{1}Dv_{2}\,\varphi\left(\sqrt{Q-q_{1}}v_{1}+\sqrt{q_{1}}x_{1}\right)\,\varphi\left(\sqrt{Q-q_{2}}v_{2}+\frac{p}{\sqrt{q_{1}}}x_{1}+\sqrt{q_{2}-\frac{p^{2}}{q_{1}}}x_{2}\right)\\ &=\int Dx_{1}Dx_{2}\,\varphi\left(\sqrt{Q}x_{1}\right)\,\varphi\left(\frac{p}{\sqrt{Q}}\,x_{1}+\sqrt{Q-\frac{p^{2}}{Q}}x_{2}\right)=\Delta_{Q}(p)\\ \end{split}$ (74)
@@
$\begin{split}\langle\langle\varphi_{r}\rangle_{v}\langle\varphi_{r}\rangle_{v}\rangle_{x}=\langle\langle\varphi_{r}\rangle_{v}^{2}\rangle_{x}=\Delta_{Q}(q_{r})\end{split}$ (75)
@@
Secondly let’s analyze the terms which contain only $\tilde{\varphi}$, i.e. $T\equiv\langle\langle\tilde{\varphi}\rangle^{2}_{v}\rangle_{x}-\langle\tilde{\varphi}\rangle_{v,x}^{2}$ and $\Xi\equiv\langle\tilde{\varphi}^{2}\rangle_{v,x}-\langle\langle\tilde{\varphi}\rangle^{2}_{v}\rangle_{x}$. The terms $\langle\tilde{\varphi}\rangle_{v,x}$ and $\langle\tilde{\varphi}^{2}\rangle_{v,x}$ are easy to analyze since the integrand involves a sum of 4 uncorrelated Gaussian variables, which is Gaussian. We therefore get
$\langle\tilde{\varphi}\rangle_{v,x}$ $=\int Dy\,\varphi(\sqrt{Q}y)=\sqrt{\Delta_{Q}(0)}\,,$ (76a)
$\langle\tilde{\varphi}^{2}\rangle_{v,x}$ $=\int Dy\,\varphi^{2}(\sqrt{Q}y)=\Delta_{Q}(Q)\,.$ (76b)
@@
$\begin{split}\langle\langle\tilde{\varphi}\rangle^{2}_{v}\rangle_{x}&=\int Dx_{1}Dx_{2}\left[\int Dv_{1}Dv_{2}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{r}+\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)\right]^{2}\\ &=\int Dx\left[\int Dy\,\varphi\left(\frac{\sqrt{Q-2Q\gamma(1-\gamma)-\gamma^{2}q_{1}-(1-\gamma)^{2}q_{2}}}{c_{\gamma}}y+\frac{\sqrt{\gamma^{2}q_{1}+(1-\gamma)^{2}q_{2}+2\gamma(1-\gamma)p}}{c_{\gamma}}\,x\right)\right]^{2}\\ &=\Delta_{Q}\left(\frac{\gamma^{2}q_{1}+(1-\gamma)^{2}q_{2}+2\gamma(1-\gamma)p}{c_{\gamma}^{2}}\right)\end{split}$ (77)
@@
The last computation concerns the correlations between the function $\varphi_{r}$ and $\tilde{\varphi}$, which appear in the variables $U_{r}\equiv\langle\langle\varphi_{r}\rangle_{v}\langle\tilde{\varphi}\rangle_{v}\rangle_{x}-\langle\varphi\rangle_{v,x}\langle\tilde{\varphi}\rangle_{v,x}$, $\Omega_{r}\equiv\langle\langle\tilde{\varphi}\varphi_{r}\rangle_{v}\rangle_{x}-\langle\langle\tilde{\varphi}\rangle_{v}\langle\varphi_{r}\rangle_{v}\rangle_{x}$. We need therefore to evaluate the following two quantities $\langle\langle\tilde{\varphi}\varphi_{r}\rangle_{v}\rangle_{x}$ and $\langle\langle\tilde{\varphi}\rangle_{v}\langle\varphi_{r}\rangle_{v}\rangle_{x}$, for $r=1,2$. We are going to analyze the case $r=1$, the other can be obtained by symmetry. We have
@@
$\begin{split}\langle\langle\tilde{\varphi}\varphi_{1}\rangle_{v}\rangle_{x}&=\int Dx_{1}Dx_{2}Dv_{1}Dv_{2}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{r}+\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)\varphi\left(\sqrt{Q-q_{1}}v_{1}+\sum_{s}\mathcal{T}_{1s}x_{s}\right)\\ &=\int Dx_{1}Dx_{2}Dv_{1}Dv_{2}\,\varphi(x_{1})\\ &\times\varphi\left(\frac{\gamma Q+(1-\gamma)p}{\sqrt{Q}c_{\gamma}}x_{1}-\frac{(1-\gamma)p\sqrt{Q-q_{1}}}{\sqrt{q_{1}Q}c_{\gamma}}v_{1}+\frac{(1-\gamma)\sqrt{Q-q_{2}}}{c_{\gamma}}v_{2}+\frac{1-\gamma}{c_{\gamma}}\sqrt{q_{2}-\frac{p^{2}}{q_{1}}}x_{2}\right)\\ &=\int Dx_{1}Dx_{2}\,\varphi(\sqrt{Q}x_{1})\varphi\left(\frac{\gamma Q+(1-\gamma)p}{\sqrt{Q}c_{\gamma}}x_{1}+\frac{1-\gamma}{c_{\gamma}}\sqrt{Q-\frac{p^{2}}{Q}}x_{2}\right)=\Delta_{Q}\left(\frac{\gamma Q+(1-\gamma)p}{c_{\gamma}}\right)\end{split}$ (78)
@@
Notice that this does not depend on either $q_{1}$ nor $q_{2}$. The case $r=2$ can be obtained by sending $\gamma\to 1-\gamma$. Let’s analyze the term $\langle\langle\tilde{\varphi}\rangle_{v}\langle\varphi_{r}\rangle_{v}\rangle_{x}$ in the $r=1$ case
@@
$\begin{split}\langle\langle\tilde{\varphi}\rangle_{v}\langle\varphi_{1}\rangle_{v}\rangle_{x}&=\int Dx_{1}Dx_{2}Dv_{1}Dv_{2}Dv_{3}\,\varphi\left(\sqrt{Q-q_{1}}v_{3}+\sum_{s}\mathcal{T}_{1s}x_{s}\right)\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{r}+\sum_{s}\mathcal{T}_{rs}x_{s}}{c_{\gamma}}\right)\right)\\ &=\int Dx_{1}Dx_{2}\,\varphi(\sqrt{Q}x_{1})\varphi\left(\frac{\gamma q_{1}+(1-\gamma)p}{\sqrt{Q}c_{\gamma}}x_{1}+\sqrt{Q-\frac{(\gamma q_{1}+(1-\gamma)p)^{2}}{Qc_{\gamma}^{2}}}x_{2}\right)\\ &=\Delta_{Q}\left(\frac{\gamma q_{1}+(1-\gamma)p}{c_{\gamma}}\right)\end{split}$ (79)
@@
$\begin{split}\langle\langle\tilde{\varphi}\rangle_{v}\langle\varphi_{2}\rangle_{v}\rangle_{x}=\Delta_{Q}\left(\frac{(1-\gamma)q_{2}+\gamma p}{c_{\gamma}}\right)\end{split}$ (80)